OpenStudy (candy13106):

If a bus left the station at 8:21 a.m. and arrived at its destination at 10:51 p.m. on the next day, how long did the trip take?

OpenStudy (anonymous):

31.2 hours

OpenStudy (candy13106):

thats no the choices

OpenStudy (candy13106):

37 hours and 30 minutes 38 hours and 30 minutes 39 hours and 30 minutes 26 hours and 30 minutes

OpenStudy (anonymous):

Wat Are The Choices?

OpenStudy (candy13106):

up

OpenStudy (candy13106):

hello

OpenStudy (anonymous):

I will Have To Say 26 hours and 30 minutes is that right?

OpenStudy (candy13106):

its wrong

OpenStudy (anonymous):

@hartnn Can You Help With This Problem?

hartnn (hartnn):

yeah, 8:21 am to 8.21am next day will be 24 hrs 8.21 am to 10.51 pm will be 12+2.30 hrs so, total = 24+12+2.30 =....hrs

OpenStudy (candy13106):

oh thanks